OpenStudy (anonymous):

The theoretical probability of undesirable side effects resulting from taking Grebex is 1 in 11. If 121 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

1/11 = x /121 x = 11 11 people

OpenStudy (anonymous):

@keebler jimmy is right

OpenStudy (lalaly):

EV = p(x) * x = 1/11 * 121 = 11